OpenStudy (anonymous):

sinx/1-cosx+sinx/1+cosx=2 csc x how do i prove that this is a identity?

OpenStudy (abmon98):

\[sinx/1-cosx+sinx/1+cosx\] \[sinx(1+cosx)+sinx*(1-\cos)/(1-cosx)(1+cosx)\] \[sinx+sinxcosx+sinx-sinxcosx/1+cosx-cosx-cosx^2\] \[2sinx/\sin^2x\]

OpenStudy (abmon98):

\[2*sinx/sinx*sinx\]

OpenStudy (abmon98):

1/sinx=cscx 2/sinx=2cscx

OpenStudy (abmon98):

2*sinx/sinx*sinx cancel out sinx so we are left with 2/sinx

OpenStudy (sweetnikkie072):

LHS= sin x^2/(1-cos x) + sin x/(1+cos x) =take LCM we get, {sin x(1+cos x)+sin x (1- cos x)} / (1-cos^2 x) cancel +cos x and -cos x as it = 0 we are left with {sin x (1+1)} / sin^2 x => 2 sin x/sin^2 x cancel sin x from numerator and denominator we get, 2 (1/sin x) now taking a look at RHS 2 cosec x 2(1/sinx) (as cosec is reciprocal of sin) now LHS=RHS hence proven .. HOPE U CAN UNDERSTAND THE STEPS..I TRIED MY BEST!